Bob and Carol want to hire Abel Refinishing to sand and refinish the dining room floor to match the floor in the living room. Abel charges $2.89 per square foot to sand and refinish a hardwood floor. The dining room is rectangular and measures 17 feet 6 inches by 11 feet 6 inches. Find the area of the dining room floor, rounded up to the next square foot, and the cost of the work.

Answer by fractalier(6550)

Area is length times width here, so
A = 17 1/2 feet x 11 1/2 feet
A = 35/2 x 23/2
A = 805/4 = 201.25 square feet
rounded up to 202...then multiply this by $2.89 and get the cost...
$583.78